The camera control computer is down due to the storm. It should be back up tomorrow.
Category Archives: Maintenance
cloud limit raised to 70%
Cloud limit was raised to 70%. This should allow some backlogged requests to get through.
Weather station back, RCT working
The MAWS is back and working fine. The serial port the MAWS was connected to showed signs of intermittent failure. A backup card was installed, and the port changed (ttyS0 -> ttyS4). The opportunity was taken to re-evacuate the dewar, which was not cooling completely. The telescope is on and functioning.
RCT weather station inoperative
The vaisala weather station is not responding. The telescope will be down for at least a few days until the station can be inspected.
The focus was off last night. I’ll interrupt the queue at the beginning of tonight to re-calibrate it.
New UPS ordered
The new building UPS has been ordered and is en route.
A few servicing issues resolved
Brevin went up last night (6/22) to do a quick servicing on the mirror support compressor, the axis video system, and the oculus camera control computer. In short, all issues are fixed, and the telescope is back to normal operations. Please check images from tonight (6/23) to insure that the telescope is functioning properly. Please reply to this message, or email me personally, if you still find issues. Details on each issue are below.
Camera control machine down
The camera control computer is not responding following a loss in power on July 20. Brevin has been alerted.
RCT experienced a power out at 00:10 local time. Telescope was restarted and homed at 11:00 mountain.
Several images the last few nights are blank. I suspect the dome encoders are off, and have re-homed the encoder. Please send comments if your images are still blank.